9. H ? ow can a bond have a negative rate of return A) if the current yield is greater than the coupon rate B) if the current yield is less than the coupon rate C) if the rate of capital gains is less than the current yield D) if the rate of capital loss exceeds the current yield 10. What is the rate of return on a bond with a coupon of $55 that was purchased for $900 and sold one year later for $950? A) 5.56% B) 6.11% C) 11.67% D) 12.43% 11. If the real interest rate is 1.4% and the nominal interest rate is 0.6%, expected inflation equals A)-2% B) -0.8% C) 0.8% D) 2%
Answer #1

9. How can a bond have a negative rate of return?
A bond will have negative rate of the return if the capital loss exceeds the current yield.
Option D. if the rate of capital loss exceeds the current yield

10. Rate of return = (capital gain + coupon payment) / purchase price = (950-900 + 55 ) / 900 = 11.67%.
Option C. 11.67%

11. Using the Fisher equation, real interest rate ≈ nominal interest rate − inflation rate
Thus, expected inflation rate = 0.6 - 1.4 = -0.8%
Option B. -0.8%

Hope this helped!